A plant operation has fixed costs of $2,000,000

per year, and its output capacity is 100,000 electrical

appliances per year. The variable cost is $40 per unit, and

the product sells for $90 per unit.

a. Construct the economic breakeven chart.

b. Compare annual profit when the plant is operating

at 90% of capacity with the plant operation at 100%

capacity. Assume that the first 90% of capacity output

is sold at $90 per unit and that the remaining 10% of

production is sold at $70 per unit.
Answer #1

a. Break even point in terms of units = Fixed costs/ ( Selling price - Variable cost) = 2,000,000 / ( 90 -40) = 40,000 units

b. Break even point in terms of dollars = 40,000 * 90 = 3,600,000

c. Since nothing is explicitly mentioned, it is assumed that 100 % of the capacity has been used.

d. New break even point = Fixed costs/ ( Selling price - Variable cost) = 2,000,000 / ( 90 -55) = 57143 units

% reduction in profit = (3,000,000 - 1,500,000) / 3,000,000 = 50%


Profit at variable cost of $40Profit at variable cost of $55
Sales                9,000,000                                     9,000,000
Less: Variable cost                4,000,000                                     5,500,000
Contribution margin                5,000,000                                     3,500,000
Less: Fixed costs2,000,0002,000,000
Profit                3,000,000                                     1,500,000


answered by: anonymous